OpenStudy (anonymous):

When Erik went to bed one winter night in northern Minnesota, the temperature was -2°F. By the time he got up the next morning, the temperature had dropped to -17°F. Assuming the temperature dropped steadily through the night, what was the temperature in the middle of the night?

OpenStudy (anonymous):

It depends upon when he went to bed. Are you people STILL using Fahrenheit???

OpenStudy (anonymous):

Idk.. I got it from a book & I didn' t understand it so I put it up here for help

OpenStudy (rogue):

When did he go to bed?

OpenStudy (rogue):

And when is "morning?"

OpenStudy (anonymous):

In Europe the question would be what does the degrees F mean?

OpenStudy (anonymous):

This doesn't help me at all :[ The question was worded poorly so maybe thats why I don't get it, Can anyone help me on the question?

OpenStudy (anonymous):

I think it's too simple. Suppose he goes to bed at 10pm and gets up at 8am and the temperature falls by 15 degrees, what was the temperature halfway through the night (at 3am)?

OpenStudy (rogue):

Alright, lets assume the kid went to bed at 10 P.M. and he woke up at 6 A.M. There is an 8 hour interval in between. In 8 hours, the temperature dropped 15 degrees from -2 to -17. If the temperature dropped at a steady rate, it dropped 15/8 = 1.875 degrees per hour. From 10 P.M., to midnight, there is a 2 hour interval. 2*1.875 = 3.75. So -2 - 3.75 = -5.75 ---> -6 degrees F.

OpenStudy (rogue):

By axis's assumption, the drop rate would be 1.5 deg/hour. So 10 p.m. to midnight, it would drop 3 degrees. -2 deg ---> -5 deg.

OpenStudy (rogue):

Its a dumb question, don't worry about it too much.

OpenStudy (anonymous):

It depends what you define the middle of the night to be!!!

OpenStudy (rogue):

Middle of the night = midnight.

OpenStudy (rogue):

ohhhh, nooo, we're over complicating it.

OpenStudy (anonymous):

Well they gave me answers to choose from, it says: -19°F -9.5°F -15°F -8.5°F I just don't know how to get the answer

OpenStudy (rogue):

(-2 - 17)/ 2 = -9.5 deg. F

OpenStudy (anonymous):

-8.5

OpenStudy (anonymous):

o.o which is it... -8.5 or -9.5 ?

OpenStudy (rogue):

-9.5 :P

OpenStudy (anonymous):

They are saying it falls from -2 by half of 15 degrees = -8.5

OpenStudy (rogue):

half of 15 = 7.5 -2 - 7.5 = 9.5

OpenStudy (rogue):

-9.5*

OpenStudy (anonymous):

My mistake - sorry!

OpenStudy (rogue):

Hehe :P dw, we all make mistakes ;D

OpenStudy (anonymous):

Ty :D

OpenStudy (rogue):

yw :)

OpenStudy (rogue):

Can you guys give me some motivvation to stop procrastianting my homework?

OpenStudy (rogue):

I'm such a lazy bum :/

OpenStudy (anonymous):

Get Perspective! :D Get an Attitude! Get Your Eye on the Prize! You can do this! After u finish your homework you can play =D

OpenStudy (anonymous):

& You won t have to worry about it anymore <3

OpenStudy (anonymous):

I'm off then. Bye.

OpenStudy (anonymous):

byee ty :]

OpenStudy (rogue):

Haha, thanks for that 'drilling.' I'm still buried in work to do and I only have 2 days left, gonna have to pull some allnighters.

OpenStudy (anonymous):

Heh thats what i'm doing now. Pretty sure u have alot more than I do. Good luck on them :]

OpenStudy (rogue):

Thanks, fell free to ask me if you need help :)

OpenStudy (anonymous):

Theres one I don't get :o It says Find the midpoint between two points on a number line if one of the points is at -7, and the other point is at 12. How do I do that o:

OpenStudy (rogue):

Add the two values up and divide by 2. Pretty much take the average.

OpenStudy (anonymous):

Ohh ok :o

OpenStudy (anonymous):

Thanks I got 2.5 :D

OpenStudy (rogue):

good job =)

OpenStudy (anonymous):

Thanks =]